OpenStudy (anonymous):

What two Numbers can you add to get 52 and subtract and get 22?

OpenStudy (hba):

Let the two numbers be x and y x+y=52 x-y=22 Now solve for x and y :)

OpenStudy (slaaibak):

x + y = 52 x - y = 22 adding both equations: 2x = 74 x = 37 substituting 37 in the first equation in x: 37 + y = 52 y = 15 testing: 37 + 15 = 52 37 - 15 = 22 it works. so it's 37 and 15.